# Hybrid LLM-DDQN based Joint Optimization of V2I Communication and Autonomous Driving

### v2

@@ III-B Example Design and Distance-based Selection
We select the top-$K$ examples with the closest Euclidean distance. After accumulating good and bad examples, we need to prioritize the relevant appropriate examples. In this work, we select the top-$K$ relevant examples per category (good or bad) via Euclidean distance before forwarding them to LLM agents. Considering we want to filter the relevant good examples from $\mathcal{E}^{\mathrm{good}}_{t-1}$ for the current state $\mathcal{S}^{\mathrm{AD}}_{t}$, which dimension is $M_{1}\times N_{f}$. We flat this 2D state to a scalar series $\{\left(\mathbf{x}^{1}_{t},\mathbf{y}^{1}_{t},\mathbf{v}^{1}_{t},\boldsymbol{\psi}^{1}_{t}\right),\dots,\left(\mathbf{x}^{M_{1}}_{t},\mathbf{y}^{M_{1}}_{t},\mathbf{v}^{M_{1}}_{t},\boldsymbol{\psi}^{M_{1}}_{t}\right)\}$, euclidean distance $d$ is defined as:
$d=\sum_{j=1}^{M_{1}}\left\lVert\left(\mathbf{x}^{j}_{t},\mathbf{y}^{j}_{t},\mathbf{v}^{j}_{t},\boldsymbol{\psi}^{j}_{t}\right)-\left(\mathbf{x}^{j}_{E_{t}},\mathbf{y}^{j}_{E_{t}},\mathbf{v}^{j}_{E_{t}},\boldsymbol{\psi}^{j}_{E_{t}}\right)\right\rVert,$

### v3

@@ III-B Example Design and Distance-based Selection
We select the top-$K$ examples with the closest Euclidean distance. After accumulating good and bad examples, we need to prioritize the relevant appropriate examples. In this work, we select the top-$K$ relevant examples per category (good or bad) via Euclidean distance before forwarding them to LLM agents. Considering we want to filter the relevant good examples from $\mathcal{E}^{\mathrm{good}}_{t-1}$ for the current state $\mathcal{S}^{\mathrm{AD}}_{t}$, which dimension is $M_{1}\times N_{f}$. We flat this 2D state to a scalar series $\{\left(\mathbf{x}^{1}_{t},\mathbf{y}^{1}_{t},\mathbf{v}^{1}_{t},\boldsymbol{\psi}^{1}_{t}\right),\dots,\left(\mathbf{x}^{M_{1}}_{t},\mathbf{y}^{M_{1}}_{t},\mathbf{v}^{M_{1}}_{t},\boldsymbol{\psi}^{M_{1}}_{t}\right)\}$, euclidean distance $d$ is defined as: $d=\sum_{j=1}^{M_{1}}\left\lVert\left(\mathbf{x}^{j}_{t},\mathbf{y}^{j}_{t},\mathbf{v}^{j}_{t},\boldsymbol{\psi}^{j}_{t}\right)-\left(\mathbf{x}^{j}_{E_{t}},\mathbf{y}^{j}_{E_{t}},\mathbf{v}^{j}_{E_{t}},\boldsymbol{\psi}^{j}_{E_{t}}\right)\right\rVert,$